# NAVINACT: Combining Navigation and Imitation Learning for Bootstrapping Reinforcement Learning

### v1

@@ Abstract
Reinforcement Learning (RL) has shown remarkable progress in simulation environments, yet its application to real-world robotic tasks remains limited due to challenges in exploration and generalization. To address these issues, we introduce NAVINACT, a framework that chooses when the robot should use classical motion planning based navigation and when it should learn a policy. To further improve the efficiency in exploration, we use imitation data to bootstrap the exploration. NAVINACT dynamically switches between two modes of operation: navigating to a waypoint using classical techniques when away from the objects and reinforcement learning for fine-grained manipulation control when about to interact with objects. NAVINACT consists of a multi-head architecture composed of ModeNet for mode classification, NavNet for waypoint prediction, and InteractNet for precise manipulation. By combining the strengths of RL and Imitation Learning (IL), NAVINACT improves sample efficiency and mitigates distribution shift, ensuring robust task execution. We evaluate our approach across multiple challenging simulation environments and real-world tasks, demonstrating superior performance in terms of adaptability, efficiency, and generalization compared to existing methods.In both simulated and real-world settings, NAVINACT demonstrates robust performance. In simulations, NAVINACT surpasses baseline methods by 10-15% in training success rates at 30k samples and by 30-40% during evaluation phases. In real-world scenarios, it demonstrates a 30-40% higher success rate on simpler tasks compared to baselines and uniquely succeeds in complex, two-stage manipulation tasks. Datasets and supplementary materials can be found on our website.
@@ 1 Introduction
Figure 1: Multi-headed architecture of NAVINACT: During training, NAVINACT learns to predict waypoints, low-level actions, and the operational mode at each time step. One network (InteractNet) predicts the low-level action $a_{t}$ and the other network (ModeNet) predicts mode $m_{t}$. A separate network (NavNet) predicts the high-level waypoint $w_{t}$. At test time, the system samples $m_{t}$ and either navigates to a waypoint (when $m_{t}=0$) using the predicted waypoint or follows a dense action (when $m_{t}=1$). The architecture allows for dynamic switching between navigation and interaction modes, facilitating seamless transitions and robust performance in complex tasks. An example of how navigation and interaction modes are integrated during execution is shown on the right.
@@
To enhance sample efficiency and generalizability, researchers have explored imitation learning (IL) as a means to accelerate learning by leveraging demonstration data. In hierarchical frameworks like LAVA [5], IL is used to parameterize low-level policies based on high-level visual inputs, enabling the system to effectively handle varying types and positions of food during robotic assisted feeding tasks. However, IL alone can lead to distributional shift issues when applied outside the training context. Our focus in this paper is to improve the sample complexity and generalizability of robot learning specifically for manipulation tasks. Our key idea is to improve the exploration process relying on classical motion planning algorithms and imitation data. Specifically, we use motion planning when the robot is “away” from objects and applying RL only when the robot is “near” objects during interactions that require fine control. This strategy reduces the complexity of the learning problem and enhances sample efficiency. Further improvements in sample efficiency can be achieved through imitation learning (IL), though IL alone can raise generalizability issues, which we address by bootstrapping IL with RL.
Our novel approach leverages the hierarchical nature of robotic tasks by implementing a dynamic policy framework that switches between navigating to high-level waypoints and engaging in fine-grained manipulation. We build on HYDRA [29] which introduced the idea of switching between sparse (classical motion planning) and dense (learning) modes. However, the dense mode in HYDRA uses IL which is prone to distributional shift near objects, our method integrates IL with RL, mitigating these shifts and improving task robustness.
@@
We introduce NAVINACT (NAVigation and INterACTion), a framework that integrates navigation and imitation learning to bootstrap reinforcement learning. Our approach comprises three key components: NavNet for strategic waypoint planning based on visual inputs; ModeNet that dynamically switches between navigation and interaction modes; and InteractNet that performs fine-grained manipulation tasks with precision as shown in Fig. 5. The key contributions of this paper are:
  • We present NAVINACT, a hierarchical policy framework designed to integrate navigation and interaction for complex robotic manipulation tasks.
@@
  • We thoroughly tested NAVINACT in both simulated and real-world settings using a UR3e robot equipped with a Robotiq Hand-e gripper and Realsense cameras. The framework achieved impressive success rates: 85±5% during simulation training and 80±5% in evaluations with unseen environments. In real-world tests, NAVINACT achieved up to 90% success on simpler tasks and 75% on complex two-stage tasks within 30k samples, highlighting its efficiency and adaptability.
@@ 2.1 Incorporating Model Priors and Refining Action Spaces
Research has increasingly concentrated on embedding structural priors within models to stabilize performance across variable conditions. Object-centric representations and pretrained state models from diverse task datasets have proven to enhance generalization and efficiency [1, 2, 3]. Approaches like [29] HYDRA have refined action spaces to align with these insights. However, NAVINACT goes a step further by dynamically switching action modes, integrating both high-level strategic planning and precise low-level task executions. This not only enhances navigation and manipulation capabilities in intricate settings but also addresses the challenges of generalization and extensive data requirements found in temporal action abstractions and parameterized motion primitives [4, 7, 8, 9, 11, 12]. Building on the foundations laid by previous models, NAVINACT effectively minimizes errors and advances the capabilities of action representation in complex environments, making it a significant advancement in both model integration and action space refinement.
@@ 2.2 Planning-Driven Reinforcement Learning
Integrating motion planning with RL has been explored to combine the benefits of both paradigms [13, 14, 15, 16, 17, 18, 19]. For example, GUAPO [13] focuses on single-stage tasks, maintaining RL agents within low pose-estimator uncertainty areas, while Plan-Seq-Learn decomposes complex tasks into sequences of subtasks guided by large language models (LLMs). However, these methods often come with high inference costs or inefficiencies in task decomposition. Our NAVINACT method contrasts these approaches by offering an end-to-end learning solution that is not only faster but also more adaptable to real-time applications, with the potential for further enhancement through LLM integration.
@@ 2.3 Reinforcement Learning with Prior Demonstrations
Sample-efficient RL often struggles in sparse reward environments, where initial learning cues are vital. Integrating prior data or human demonstrations into the replay buffer—an approach employed in methods like Reinforcement Learning from Prior Data (RLPD)[31] and IBRL [28]—has proven effective. These methods enhance RL algorithms by oversampling demonstrations during training, significantly improving performance in continuous control domains with techniques like normalization and Q-ensembling [2, 28]. Our NAVINACT framework builds on these principles by dynamically adjusting its strategy based on environmental cues and task demands, using a multi-headed architecture that optimizes both navigation and interaction. This approach not only leverages prior demonstrations for improved initial guidance but also adapts in real-time to enhance task execution and generalization in complex robotic manipulation tasks.
@@ 2.4 Advancing RL Sample Efficiency
Recent advancements in RL have prioritized enhancing sample efficiency through innovative regularization strategies. Techniques like RED-Q [20] and DropoutQ [24] apply regularization to the Q-function, utilizing ensembling or dropout to achieve higher update-to-data (UTD) ratios and expedited convergence. These approaches prove particularly effective in state-based RL, while in pixel-based settings, strategies such as image augmentation with random shifts [25] boost performance without additional computational demands. Building upon these concepts, the IBRL [28] framework integrates imitation learning (IL) with RL, harnessing a pre-trained IL policy to direct the RL training, enhancing the initial learning curve. However, IBRL still necessitates frequent updates to achieve optimal performance levels. In contrast, our NAVINACT method refines this approach by significantly reducing the number of updates required, effectively merging the benefits of IL, RL as well as mode switching and navigation to heighten sample efficiency and streamline learning processes.
3 NAVINACT
@@
Figure 2: ModeNet architecture for classifying modes (navigate vs. interact) based on visual inputs.
Figure 3: NavNet architecture for predicting waypoints to guide high-level navigation tasks.
Our proposed method, NAVINACT, features a multi-headed architecture designed for efficient robotic manipulation. The system is composed of three main components: ModeNet, NavNet, and InteractNet, each contributing to different levels of task execution. The overall architecture, as illustrated in Fig. 1, includes heads $\pi^{M}_{\theta}:\mathcal{S}\rightarrow\{0,1\}$, $\pi^{W}_{\theta}:\mathcal{S}\rightarrow P$, and $\pi^{A}_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$, corresponding to mode classification, high-level waypoint prediction, and low-level action prediction respectively, where $P$ is robot proprioceptive data such as end effector poses and gripper widths. This design enables dynamic switching between coarse and fine-grained control during task execution.
@@ 3.1 ModeNet: A Vision-Based Mode Classification Network
ModeNet is a deep learning architecture that classifies the operational mode of the robot based on visual input. It determines whether the robot should operate in navigation mode, or in interaction mode, as shown in Fig. 5.
ModeNet utilizes a ResNet backbone as shown in Fig. 3. The architecture consists of a feature extractor followed by a mode classifier. The feature extractor employs a pretrained ResNet model to capture high-level spatial and contextual features from input images, while the mode classifier processes these features to predict the robot’s operational mode.
@@ 3.2 NavNet: A Vision-Based Waypoint Prediction Network
NavNet is designed to predict waypoints for robotic manipulation tasks by processing both visual and proprioceptive inputs. It generates high-level action outputs that guide the robot towards the target objects in the environment. NavNet combines a convolutional neural network (CNN) for visual processing with fully connected layers that integrate visual features with proprioceptive data, Fig. 3. This design allows for efficient waypoint prediction, facilitating robust navigation in complex environments. Once we have the predicted waypoint, we use a sampling-based motion planner, AIT* [39], to perform motion planning.
Algorithm 1 NAVINACT. Major modifications with framework highlighted in blue.
@@
3:  Train ModeNet on dataset $D_{\text{mode}}$ to classify mode (navigate vs interact).
4:  Train NavNet on dataset $D_{\text{nav}}$ to predict waypoints for navigation tasks.
@@
12:        Generate navigation action $a^{\text{Nav}}_{t}$ based on waypoint $w_{t}$
13:        Execute navigation action $a^{\text{Nav}}_{t}$
@@ 3.3 InteractNet: Low-Level Action Prediction
Fig. 4 illustrates how often IL actions are chosen throughout training. Early on, the predominance of RL actions is noticeable due to initial high “fake” Q-values. As the system updates and learns from real performance data, IL actions become more frequent, underscoring their effectiveness. Over time, while the reliance on IL decreases as the RL policy matures, IL actions remain crucial, particularly in complex tasks where RL alone is insufficient for full policy convergence. This ongoing use of IL actions at all stages indicates their integral role in ensuring robust and efficient decision-making within NAVINACT.
@@ 4 Experiments
In this section, we present a comprehensive evaluation of NAVINACT, focusing on its sample efficiency, generalizability, and performance in real-world scenarios. We conducted experiments in three challenging simulation environments (Fig. 8) and two complex real-world tasks (Fig. 10).
@@ 4.1 Training and evaluation of ModeNet and NavNet
Figure 6: This figure displays the evaluation setups for three MetaWorld tasks: Assembly, BoxClose, and CoffeePush, with randomized initial gripper positions. Each subfigure illustrates how the gripper’s starting position varies in these environments, testing NAVINACT’s ability to adapt to different initial conditions.
@@
ModeNet is trained on a dataset comprising 1,500 environmental camera images for simulation. The ground truth labels were determined by a distance threshold between the end-effector and the object, effectively deciding whether the robot should be in a navigation mode (moving towards an object) or an interaction mode (engaging directly with an object). Data augmentation techniques like rotation and contrast adjustment were applied during training. The network utilizes cross-entropy loss and is optimized with Adam. The network’s performance is evaluated using metrics such as accuracy, precision, recall, and F1 score. ModeNet achieved an accuracy score of 0.89, an F1 score of 0.85, a recall score of 0.9, and a precision score of 0.8, demonstrating its effectiveness in accurately classifying operational modes for robotic tasks.
@@ 4.1.1 NavNet
NavNet is trained on the same dataset of 1,500 environmental camera images as ModeNet. The ground truth for NavNet involves waypoints defined as positions 2 cm above the objects, guiding the robot on where to navigate next. Like ModeNet, data augmentation techniques such as rotation and contrast adjustments were used to enhance training robustness. NavNet employs Mean Squared Error (MSE) loss and is optimized with the Adam optimizer. It achieved an impressive accuracy score of 0.93, indicating high effectiveness in predicting accurate navigation waypoints based on visual inputs. Fig. 7 shows basic demo of NAVINACT in navigation mode.
4.1.2 ModeNet and NavNet for realworld experiments
For real-world experiments, ModeNet and NavNet were trained using a more extensive dataset, incorporating both wrist and environment camera images totaling 6,200. This approach enhanced the diversity and representativeness of the training data, ensuring robustness and accuracy in real-world applications.

### v2

# PLANRL: A Motion Planning and Imitation Learning Framework to Bootstrap Reinforcement Learning

@@ Abstract
Reinforcement Learning (RL) has shown remarkable progress in simulation environments, yet its application to real-world robotic tasks remains limited due to challenges in exploration and generalization. To address these issues, we introduce PLANRL, a framework that chooses when the robot should use classical motion planning and when it should learn a policy. To further improve the efficiency in exploration, we use imitation data to bootstrap the exploration. PLANRL dynamically switches between two modes of operation: reaching a waypoint using classical techniques when away from the objects and reinforcement learning for fine-grained manipulation control when about to interact with objects. PLANRL architecture is composed of ModeNet for mode classification, NavNet for waypoint prediction, and InteractNet for precise manipulation. By combining the strengths of RL and Imitation Learning (IL), PLANRL improves sample efficiency and mitigates distribution shift, ensuring robust task execution. We evaluate our approach across multiple challenging simulation environments and real-world tasks, demonstrating superior performance in terms of adaptability, efficiency, and generalization compared to existing methods. In simulations, PLANRL surpasses baseline methods by 10-15% in training success rates at 30k samples and by 30-40% during evaluation phases. In real-world scenarios, it demonstrates a 30-40% higher success rate on simpler tasks compared to baselines and uniquely succeeds in complex, two-stage manipulation tasks. Datasets and supplementary materials can be found on our website.
@@ 1 Introduction
Figure 1: Architecture of PLANRL: During training, PLANRL learns to predict waypoints, low-level actions, and the operational mode at each time step. One network (InteractNet) predicts the low-level action $a_{t}$ and the other network (ModeNet) predicts mode $m_{t}$. A separate network (NavNet) predicts the high-level waypoint $w_{t}$. At test time, the system samples $m_{t}$ and either moves to a waypoint (when $m_{t}=0$) using the predicted waypoint or follows a dense action (when $m_{t}=1$). The architecture allows for dynamic switching between motion-planning and interaction modes, facilitating robust performance in complex tasks. An example of how motion planning and interaction modes are integrated during execution is shown on the right.
@@
To enhance sample efficiency and generalizability, researchers have explored imitation learning (IL) as a means to accelerate learning by leveraging demonstration data. In hierarchical frameworks like LAVA [5], IL is used to parameterize low-level policies based on high-level visual inputs, enabling the system to effectively handle varying types and positions of food during robotic-assisted feeding tasks. However, IL alone can lead to distributional shift issues when applied outside the training context. Our focus in this paper is to improve the sample complexity and generalizability of robot learning specifically for manipulation tasks. Our key idea is to improve the exploration process relying on classical motion planning algorithms and imitation data. Specifically, we use motion planning when the robot is “away” from objects and apply RL only when the robot is “near” objects during interactions that require fine control. This strategy reduces the complexity of the learning problem and enhances sample efficiency. Further improvements in sample efficiency can be achieved through imitation learning (IL), though IL alone can raise generalizability issues, which we address by bootstrapping IL with RL.
Our novel approach leverages the hierarchical nature of robotic tasks by implementing a dynamic policy framework that switches between moving to high-level waypoints and engaging in fine-grained manipulation. We build on HYDRA [29] which introduced the idea of switching between sparse (classical motion planning) and dense (learning) modes. However, the dense mode in HYDRA uses IL which is prone to distributional shift near objects, our method integrates IL with RL, mitigating these shifts and improving task robustness.
@@
We introduce PLANRL (NAVigation and INterACTion), a framework that integrates motion-planning and imitation learning to bootstrap reinforcement learning. Our approach comprises three key components: NavNet for strategic waypoint planning based on visual inputs; ModeNet that dynamically switches between motion-planning and interaction modes; and InteractNet that performs fine-grained manipulation tasks with precision as shown in Fig. 5. The key contributions of this paper are:
  • We present PLANRL, a hierarchical policy framework designed to integrate motion-planning and interaction for complex robotic manipulation tasks.
@@
  • We thoroughly tested PLANRL in both simulated and real-world settings using a UR3e robot equipped with a Robotiq Hand-e gripper and Realsense cameras. The framework achieved impressive success rates: 85±5% during simulation training and 80±5% in evaluations with unseen environments. In real-world tests, PLANRL achieved up to 90% success on simpler tasks and 75% on complex two-stage tasks within 30k samples, highlighting its efficiency and adaptability.
@@ 2.1 Incorporating Model Priors and Refining Action Spaces
Research has increasingly concentrated on embedding structural priors within models to stabilize performance across variable conditions. Object-centric representations and pretrained state models from diverse task datasets have proven to enhance generalization and efficiency [1, 2, 3]. Approaches like [29] HYDRA have refined action spaces to align with these insights. However, PLANRL goes a step further by dynamically switching action modes, integrating both high-level strategic planning and precise low-level task executions. This not only enhances motion-planning and manipulation capabilities in intricate settings but also addresses the challenges of generalization and extensive data requirements found in temporal action abstractions and parameterized motion primitives [4, 7, 8, 9, 11, 12]. Building on the foundations laid by previous models, PLANRL effectively minimizes errors and advances the capabilities of action representation in complex environments, making it a significant advancement in both model integration and action space refinement.
@@ 2.2 Planning-Driven Reinforcement Learning
Integrating motion planning with RL has been explored to combine the benefits of both paradigms [13, 14, 15, 16, 17, 18, 19]. For example, GUAPO [13] focuses on single-stage tasks, maintaining RL agents within low pose-estimator uncertainty areas, while Plan-Seq-Learn [30] decomposes complex tasks into sequences of subtasks guided by large language models (LLMs). However, these methods often come with high inference costs or inefficiencies in task decomposition. Our PLANRL method contrasts these approaches by offering an end-to-end learning solution that is not only faster but also more adaptable to real-time applications, with the potential for further enhancement through LLM integration.
@@ 2.3 Reinforcement Learning with Prior Demonstrations
Sample-efficient RL often struggles in sparse reward environments, where initial learning cues are vital. Integrating prior data or human demonstrations into the replay buffer—an approach employed in methods like Reinforcement Learning from Prior Data (RLPD)[31] and IBRL [28]—has proven effective. These methods enhance RL algorithms by oversampling demonstrations during training, significantly improving performance in continuous control domains with techniques like normalization and Q-ensembling [2, 28]. Our PLANRL framework builds on these principles by dynamically adjusting its strategy based on environmental cues and task demands, using an architecture that optimizes both motion-planning and interaction. This approach not only leverages prior demonstrations for improved initial guidance but also adapts in real-time to enhance task execution and generalization in complex robotic manipulation tasks.
@@ 2.4 Advancing RL Sample Efficiency
Recent advancements in RL have prioritized enhancing sample efficiency through innovative regularization strategies. Techniques like RED-Q [20] and DropoutQ [24] apply regularization to the Q-function, utilizing ensembling or dropout to achieve higher update-to-data (UTD) ratios and expedited convergence. These approaches prove particularly effective in state-based RL, while in pixel-based settings, strategies such as image augmentation with random shifts [25] boost performance without additional computational demands. Building upon these concepts, the IBRL [28] framework combines imitation learning (IL) with reinforcement learning (RL) by leveraging a pre-trained IL policy to guide RL training from the start, allowing the RL agent to benefit from high-quality actions provided by the IL policy. This integration significantly accelerates both exploration and training efficiency. However, IBRL still necessitates frequent updates to achieve optimal performance levels. In contrast, our PLANRL method refines this approach by significantly reducing the number of updates required, effectively merging the benefits of IL, RL as well as mode switching and motion-planning to heighten sample efficiency and streamline learning processes.
3 PLANRL
@@
Figure 2: ModeNet architecture designed to classify modes (motion-planning vs. interact) based on visual inputs.
Figure 3: NavNet architecture for predicting waypoints to guide high-level motion-planning tasks.
Our proposed method, PLANRL, features an architecture designed for efficient robotic manipulation. The system is composed of three main components: ModeNet, NavNet, and InteractNet, each contributing to different levels of task execution. The overall architecture, as illustrated in Fig. 1, includes heads $\pi^{M}_{\theta}:\mathcal{S}\rightarrow\{0,1\}$, $\pi^{W}_{\theta}:\mathcal{S}\rightarrow W$, and $\pi^{A}_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$, corresponding to mode classification, high-level waypoint prediction, and low-level action prediction respectively, where $W$ is robot end-effector position. This design enables dynamic switching between coarse and fine-grained control during task execution.
@@ 3.1 ModeNet: A Vision-Based Mode Classification Network
ModeNet is a deep learning architecture that classifies the operational mode of the robot based on visual input. It determines whether the robot should operate in motion-planning mode, or in interaction mode, as shown in Fig. 5.
ModeNet utilizes a ResNet backbone as shown in Fig. 3. The architecture consists of a feature extractor followed by a mode classifier. The feature extractor employs a pre-trained ResNet model to capture high-level spatial and contextual features from input images, while the mode classifier processes these features to predict the robot’s operational mode.
@@ 3.2 NavNet: A Vision-Based Waypoint Prediction Network
NavNet is designed to predict waypoints for robotic manipulation tasks by processing visual inputs. It generates high-level action outputs that guides the robot towards the target objects in the environment. NavNet uses a convolutional neural network (CNN) for visual processing, effectively handling the image data to produce actionable waypoints, Fig. 3. Once we have the predicted waypoint, we use a sampling-based motion planner, AIT* [39], to perform motion planning as it requires minimal setup—just collision checking—and performs well in planning tasks.
Algorithm 1 PLANRL. Major modifications with framework highlighted in blue.
@@
3:  Train ModeNet on dataset $D_{\text{mode}}$ to classify mode (motion-planning vs interaction).
4:  Train NavNet on dataset $D_{\text{nav}}$ to predict waypoints for motion-planning tasks.
@@
12:        Generate motion-planning action $a^{\text{Nav}}_{t}$ based on waypoint $w_{t}$
13:        Execute motion-planning action $a^{\text{Nav}}_{t}$
@@ 3.3 InteractNet: Low-Level Action Prediction
Fig. 4 illustrates how often IL actions are chosen throughout training. Early on, the predominance of RL actions is noticeable due to initially overestimated Q-values. As the system updates and learns from real performance data, IL actions become more frequent, recognizing their effectiveness from expert demonstrations in the replay buffer. Over time, although reliance on IL actions decreases as the RL policy matures, they remain essential in complex tasks where RL alone is insufficient for full policy convergence. This consistent integration of IL actions at all stages underscores their critical role in ensuring robust and efficient decision-making within PLANRL.
@@ 4 Experiments
In this section, we present a comprehensive evaluation of PLANRL, focusing on its sample efficiency, generalizability, and performance in real-world scenarios. We conducted experiments in three challenging simulation environments (Fig. 8) and two complex real-world tasks (Fig. 10).
@@ 4.1 Training and evaluation of ModeNet and NavNet
Figure 6: This figure displays the evaluation setups for three MetaWorld tasks: Assembly, BoxClose, and CoffeePush, with randomized initial gripper positions. Each subfigure illustrates how the gripper’s starting position varies in these environments, testing PLANRL’s ability to adapt to different initial conditions.
@@
ModeNet is trained on a dataset comprising 1,500 environmental camera images for simulation. The ground truth labels were determined by a distance threshold between the end-effector and the object, effectively deciding whether the robot should be in a motion-planning mode (moving towards an object) or an interaction mode (engaging directly with an object). Data augmentation techniques like rotation and contrast adjustment were applied during training. The network utilizes cross-entropy loss and is optimized with Adam. The network’s performance is evaluated using metrics such as accuracy, precision, recall, and F1 score. ModeNet achieved an accuracy score of 0.89, an F1 score of 0.85, a recall score of 0.9, and a precision score of 0.8, demonstrating its effectiveness in accurately classifying operational modes for robotic tasks.
@@ 4.1.1 NavNet
is trained on the same dataset of 1,500 environmental camera images as ModeNet. The ground truth for NavNet involves waypoints defined as positions 2 cm above the objects, guiding the robot on where to move next. Like ModeNet, data augmentation techniques such as rotation and contrast adjustments were used to enhance training robustness. NavNet employs Mean Squared Error (MSE) loss and is optimized with the Adam optimizer. It achieved an impressive accuracy score of 0.93, indicating high effectiveness in predicting accurate motion-planning waypoints based on visual inputs. Fig. 7 shows basic demo of PLANRL in motion-planning mode.
4.1.2 ModeNet and NavNet for real-world experiments
For real-world experiments, ModeNet and NavNet were trained using a more extensive dataset of 6200 images, incorporating both wrist and environment camera images. The dataset captures a wider range of object perspectives, lighting conditions, and spatial contexts. This expansion improves the robustness and accuracy in real-world applications.